OpenStudy (anonymous):

A twirler's baton is made of a slender metal cylinder of mass M and length L . Each end has a rubber cap of mass m, and you can accurately treat each cap as a particle in this problem. Find the total moment of inertia of the baton about the usual twirling axis (perpendicular to the baton through its center).

OpenStudy (anonymous):

moment of inertia of the metal cylinder=Im=(1/4)MR^2+(1/12)ML^2 moment of inertia of 1rubber cap=m(L/2)^2 =(mL^2)/4 since there are 2 identical caps at either ends, the moment of inertia of the caps=Ic=(mL^2)/2 Therefore, moment of inertia of the system=Im+Ic =(1/4)MR^2+(1/12)ML^2+(mL^2)/2